OpenStudy (anonymous):

Anyone good with tables??!! The following table shows the length and width of two rectangles. (The table will be in comment section) Which expression is the result of the perimeter of rectangle B minus the perimeter of rectangle A? A. 2x-4 B. 8x+10 C. 16x+20 D. 4x-8

OpenStudy (anonymous):



OpenStudy (anonymous):

Well, first you have to find the perimeter of both rectangles. Rectangle A: ((2x+8)*2)+((x-1)*2) Rectangle B: ((3x + 5)*2)+((2x-2)*2)

OpenStudy (anonymous):

Does it give you the value for x?

OpenStudy (anonymous):

Why are they squared on the end?

OpenStudy (anonymous):

They aren't that is a multiplication sign. It's hard to make a real multiplication sign on the computer, Did they give a value for x?

OpenStudy (anonymous):

oh ok no its really hard fo rme to work these out if oyu could explain some of the steps that would help alot.

OpenStudy (anonymous):

It's okay. Please answer MY question, did they give a value for x? The variable?

OpenStudy (anonymous):

No it would be up where the question is if they did

OpenStudy (anonymous):

Ugh, that's gonna make it a lot harder.

OpenStudy (anonymous):

Lol, I don't know if I can explain it to you. Sorry. Here, I will tag a few people that can help. @Luigi0210 @Destinymasha @DollyAcquah @AccessDenied @radar @MathLegend

OpenStudy (anonymous):

thank you for your time :)

OpenStudy (accessdenied):

You would find the perimeter of the rectangles while keeping the x there, then simplify like-terms. Perimeter of Rect. B = 2*(3x + 5) + 2*(2x - 2) Perimeter of Rect. A = 2*(2x + 8) + 2*(x - 1) You can expand each part out just by distributing the 2 through each factor, then add together the terms with x and the constants. As an example, I will simplify B: 2*(3x + 5) + 2*(2x - 2) 2*3x + 2*5 + 2*2x - 2*2 6x + 10 + 4x - 4 10x + 6 Then, subtract Rect. A from Rect. B:

OpenStudy (anonymous):

\[2 \times (3x + 5) + 2 \times 92x - 2)\]

OpenStudy (anonymous):

right?

OpenStudy (anonymous):

\[(6x + 10 ) + (4x - 4)\]

OpenStudy (anonymous):

10x + 6

OpenStudy (accessdenied):

Yup, that looks good for Rectangle B. Then rectangle A?

OpenStudy (anonymous):

\[2 \times (2x + 8) + 2 \times (x - 1)\]

OpenStudy (anonymous):

\[(4x + 16) + (2x - 2) \]

OpenStudy (anonymous):

6x + 14

OpenStudy (accessdenied):

Looks good. :) Then we subtract: Perimeter Rectangle B - Perimeter Rectangle A